OpenStudy (anonymous):

T: P_2 ----> R T(P(x))= int(from_1 to 1)P(x)dx find kernel of T help, please

OpenStudy (anonymous):

sorry what do you mean of p_2 ?

OpenStudy (anonymous):

it's polynomial degree 2

OpenStudy (anonymous):

@terenzreignz you have nothing to do here, it's linear algebra.

terenzreignz (terenzreignz):

sending me away? :)

OpenStudy (anonymous):

no, just not want to waste your talent.

OpenStudy (anonymous):

your strong is not this stuff, I have many question from cal need your help. but now I stuck with this question, after i have it done, I post my cal , you have no way to escape from me. hehehhe

OpenStudy (anonymous):

Who know? right? what if because of my sentence, you become strong in this field, too?

terenzreignz (terenzreignz):

Oh well...

terenzreignz (terenzreignz):

\[\huge T: \mathbb{R}_2[x]\rightarrow \mathbb{R}\] \[\huge T[f(x)]=\int\limits_{-1}^1f(x)dx\]

OpenStudy (anonymous):

hold on, P_2 not R2

terenzreignz (terenzreignz):

P_2 R_2 what's the difference we both mean polynomials of degree 2 :)

OpenStudy (anonymous):

of course, they are different, P_2 is a polynomial degree 2 and R2 is a function with 2 variables

OpenStudy (anonymous):

for example R2 is something like ax +y =0 and P2 is a0 + a1x +q2x^2

terenzreignz (terenzreignz):

don't mistake \(\large \mathbb{R}^2 \) for \(\large \mathbb{R}_2\)

OpenStudy (anonymous):

hey guy, R^2 =R_2 it's my prof terminology

terenzreignz (terenzreignz):

Well, anyway, notation aside, I mean polynomials of degree 2.

OpenStudy (anonymous):

dealt

terenzreignz (terenzreignz):

You know \[\large f(x) = ax^2 + bx + c \ \ \ \ \ a,b,c\in \mathbb{R} \ \ \ a\neq 0\]

OpenStudy (anonymous):

yeap

terenzreignz (terenzreignz):

\[\huge \int\limits_{-1}^1f(x)dx =\frac{2a}3+2c\]

OpenStudy (anonymous):

hehe...not get.

OpenStudy (anonymous):

ok, got

terenzreignz (terenzreignz):

Okay? Now \[\large \ker(T) = \left\{ax^2+bx+c \ \ | \ \ \ \frac{2a}{3}+2c = 0\right\}\]

OpenStudy (anonymous):

got it

terenzreignz (terenzreignz):

Well.... \[\large \frac{2a}{3}+2c = 0\]\[c=-\frac{a}{3}\]

OpenStudy (anonymous):

hold on, T (P(x)) = {int of (Px) | your condition, } not just ax^2 + bx +c, right?

terenzreignz (terenzreignz):

T(P(x)) = integral of P(x) that's it ker(T) that's where the conditional set kicks in

OpenStudy (anonymous):

and ker(T) when the condition =0 . I got it

OpenStudy (anonymous):

yeap we meet there

terenzreignz (terenzreignz):

Okay... so, the kernel of T is just \[\huge \left\{ ax^2 + bx - \frac{a}3 \ \ \ \ \ | \ \ \ \ \ a,b \in \mathbb{R}\right\}\]

OpenStudy (anonymous):

that's it?

terenzreignz (terenzreignz):

Yep :) Try integrating that from -1 to 1 you'll see it's zero :D

OpenStudy (anonymous):

hehehe... you think who have to say thank you to other?

OpenStudy (anonymous):

It's Terenze !!! for both black and white meanings !! cheer up

terenzreignz (terenzreignz):

I'm cheering up :)

OpenStudy (anonymous):

\[\huge T[f(x)]=\int\limits_{-1}^1f(x)dx\] Terence, write a triple int and a double int . I need those code, whatever you construct, I just need the codes

terenzreignz (terenzreignz):

to write a double integral simply write \iint here's the result of writing \iint \[\huge \iint\] To write a double integral with a region under, say, R write \iint\limits_{R} here's the result \[\Large \iint\limits_{R}\] To write a triple integral, write \iiint \[\huge \iiint\]

OpenStudy (anonymous):

how about the limit on?

OpenStudy (anonymous):

\[\huge \iint\]

terenzreignz (terenzreignz):

\iiint\limits_{} same concept basically \[\huge \iiint\limits_{S}\]

OpenStudy (anonymous):

hey, friend, I mean specific limits, some thing like int_ from 0 to 1 and then int from 1 to 2 and then int from c to d

terenzreignz (terenzreignz):

well, just write multiple integrals, then. there's no other way \int\limits_{a}^{b} \int\limits_{c}^{d} \[\huge \int\limits_a^b\int\limits_c^d\]

OpenStudy (anonymous):

thanks a lot. But I gotta run. will catch up with you later

terenzreignz (terenzreignz):

sure

OpenStudy (anonymous):

not this stuff,

OpenStudy (anonymous):

let me repost, someone messed it up .and send you the link

OpenStudy (anonymous):

@hoa there is a difference between multiple and iterated integrals.